OpenStudy (anonymous):

solve ; (14/(x-8)+2=10/(x-8)

OpenStudy (anonymous):

10

OpenStudy (anonymous):

i need work ..

OpenStudy (anonymous):

take x-8=1/z

OpenStudy (anonymous):

u can solve dat

OpenStudy (anonymous):

where do u get z?

OpenStudy (anonymous):

assuming

OpenStudy (anonymous):

assuming what?

OpenStudy (anonymous):

and at last put d value of z and u will get d result

OpenStudy (anonymous):

u just confused me.

OpenStudy (anonymous):

see x-8=1/z den eun becomes..14z+2=10z z=1/2 1/x-8=1/2 x=10

OpenStudy (anonymous):

now???

OpenStudy (anonymous):

there is no z.....

OpenStudy (anonymous):

ma chuda maderchod

OpenStudy (anonymous):

what?....